Juan rides his bike from Lawrence to Salisbury beach (30 miles in 3 hours). He divides distance (miles) by time (hours) and sees that his overall average speed is 10mph. He buys 3 Power bars and eats them. How fast (mph) must he return so that his overall average speed for the round trip (60 miles) is 15mph?THE ANSWER IS NOT 20mph.

Answer by RAY100(1637)

Velocity = Distance / Time ,,,,,,or t=d/v
.
outbound,,,,v=30/3 = 10 mph,,,,ok
.
if total distance is 60,,,and avg speed is to be 15 mph, t= 60/15 = 4 hrs
.
Since we already used 3 hrs,,,only one hr remains to go 30 miles,,,and v=d/t = 30/1 =30mph
.
Of course, neglecting any time spent at Salisbury Beach
.